OpenStudy (anonymous):

the population of the world, as of august 12, 2000, was estimated by the U.S. Census bureau to be about 5.779 billion people and growing at a rate of 1.38%. a. what was the population on august 12, 1800? b. what should the population be today, april 15, 2011? c. when will the population be 6.5 billion?

OpenStudy (anonymous):

does anyone know how to do this. i am soo confused

OpenStudy (dumbcow):

\[P _{t} =P _{0}(1+r)^{t}\] r = growth rate = .0138 For A) setup would be \[5.779=P _{0}(1.0138)^{200}\] solve for P

OpenStudy (anonymous):

do you divide the side next to p to the 5.779

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

where did the 200 come from

OpenStudy (dumbcow):

the number of years from 1800 to 2000

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you know how to do the other 2

OpenStudy (dumbcow):

B) \[P _{t}=5.779(1.0138)^{10.67}\] 10.67 is aPPROXIMATELY NUM OF years since august 2000 C) \[6.5 = 5.779(1.0138)^{t}\] solve for t using logs

OpenStudy (anonymous):

ok thanks

OpenStudy (dumbcow):

your welcome